and Y = X1. X1, X2-N(0,1) and are independent. Let Z = X² + X (a) Is Z useful to predict Y? (b) Is Y useful to predict Z?
Answer #1

Z is actually dependent upon the X1 and X2. And hence after knowing (distribution or value of) X1 and X2, we can find distribution of Z.

And as Y=X1, Z is dependent upon Y and X2. So this means after knowing Y and X2, we can get to know about Z

Z= Y2 + X22

Y2=Z-X2​​​​2

But after all Z = X1^2 + X2^2,so we can find Z using X1 and X2 and consequently Y can be found out using Z.

Hence

A) Z is useful to predict Y

B) Y is useful to predict Z